The Linux HOWTO Index

by Greg Hankins, gregh@sunsite.unc.edu v2.9.37, 25 August 1996 This document contains an index to the Linux HOWTOs and mini-HOWTOs, as well as other information about the HOWTO project.

Contents

1 What Are Linux HOWTOs? 1 2 Where Do I Get Linux HOWTOs? 1

2.1 HOWTO Translations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 HOWTO Index 3 4 HOWTO Project Developement 10

4.1 Wanted HOWTOs and mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.2 HOWTOs Currently Under Construction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3 Writing and Submitting a HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5 Copyright 12

1 What Are Linux HOWTOs?

Linux HOWTOs are documents which describe in detail a certain aspect of configuring or using Linux. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET2 HOWTO and the Printing HOWTO. HOWTOs are comprehensive docs - much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. There are several HOWTO formats available: plain text, PostScript, DVI, and HTML. In addition to the HOWTOs, there are a multitude of mini-HOWTOs. These are short, specific HOWTOs, such as the Colour-ls mini-HOWTO. They are only available in plain text format.

2 Where Do I Get Linux HOWTOs?

HOWTOs can be retrieved via anonymous FTP from the following sites:

   • ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO
   • ftp://tsx-11.mit.edu/pub/linux/docs/HOWTO
as well as many mirror sites <ftp://sunsite.unc.edu/pub/Linux/MIRRORS.html>.

You can also browse HOWTOs in HTML format <http://sunsite.unc.edu/mdw/HOWTO/> on the World

Wide Web. Many mirror sites <http://sunsite.unc.edu/mdw/hmirrors.html> also mirror the HOWTO

WWW files. sunsite.unc.edu is heavily used, so please use a mirror site if possible. HOWTOs are also posted monthly to the Usenet newsgroup comp.os.linux.answers.

2.1 HOWTO Translations

French translations of the HOWTO documents can be found at

ftp://ftp.ibp.fr/pub/linux/french/docs/HOWTO.

Japanese translations can be found at the following locations:

   • http://jf.gee.kyoto-u.ac.jp/JF/index.html (English)
   • http://jf.gee.kyoto-u.ac.jp/JF/JF.html (Japanese)
   • ftp://ftp.kuis.kyoto-u.ac.jp/Linux/JF/
   • ftp://wcarchive.cdrom.com/pub/linux/jf/

Italian translations are available at:

   • ftp://ftp.unipd.it/pub/Linux/pluto/docs/HOWTO
   • http://www.psico.unipd.it/ildp/docs/HOWTO/INDEX.html

Spanish translations are available at:

   • http://lml.ls.fi.upm.es/ jjamor/
   • ftp://luna.gui.uva.es/pub/docs-esp/
   • ftp://volcan.us.es

German translations are available at:

http://hp00.rz.tu-harburg.de/users/semb2204/install/index.html

If you know of any other translation projects, please let me know and I will add them to this list.

3 HOWTO Index

The following Linux HOWTOs are currently available: (The JE HOWTO, and SCSI HOWTO are only available in plain text format. This is because they have not been converted to the SGML source format.)

• Linux AX25 HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to configure AX25 networking for Linux. Updated 14 June 1996.

• Linux Access HOWTO, by Michael De La Rue <access-howto@ed.ac.uk>. How to use adaptive technology with Linux. Updated 1 August 1996.

• Linux BootPrompt HOWTO, by Paul Gortmaker <Paul.Gortmaker@anu.edu.au>. List of boot time arguments and overview of booting software. Updated 1 July 1996.

• Linux Bootdisk HOWTO, by Graham Chapman <grahamc@zeta.org.au>. How to create a boot/root maintenance disk for Linux. Updated 18 August 1996.

• Linux Busmouse HOWTO, by Mike Battersby <mib@deakin.edu.au>. Information on bus mouse compatibility with Linux. Updated 2 August 1994.

• Linux CDROM HOWTO, by Jeff Tranter <tranter@software.mitel.com>. Information on CD-ROM drive compatibility for Linux. Updated 10 June 1996.

• Linux Commercial HOWTO, by Martin Michlmayr <tbm@sypher.com>. Listing of commercial software products for Linux. Updated 23 August 1996.

• Linux Cyrillic HOWTO, by Alexander L. Belikoff <abel@wisdom.weizmann.ac.il>. How to configure Linux for use with the Cyrillic characterset. Updated 17 September 1995.

• Linux DNS HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to set up DNS. Updated 30 June 1996.

• Linux DOSEMU HOWTO, by Mike Deisher <deisher@dspsun.eas.asu.edu>. HOWTO about the Linux MS-DOS Emulator, DOSEMU. Updated 11 August 1995.

• Linux Danish HOWTO, by Thomas Petersen <petersen@post1.tele.dk>. How to configure Linux for use with the Danish characterset. Updated 2 June 1996.

• Linux Distribution HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. A list of Linux distributions. Updated 29 May 1996.

• Linux ELF HOWTO, by Daniel Barlow <daniel.barlow@linux.org>. How to install and migrate to the ELF binary file format. Updated 14 July 1996.

• Linux Emacspeak HOWTO, by Jim Van Zandt <jrv@vanzandt.mv.com>. How to use ’emacspeak’ with Linux. Updated 16 Aug 1996.

• Linux Ethernet HOWTO, by Paul Gortmaker <Paul.Gortmaker@anu.edu.au>. Information on Ethernet hardware compatibility for Linux. Updated 18 November 1995.

• Linux Finnish HOWTO, by Pekka Taipale <pjt@iki.fi>. How to configure Linux for use with the Finnish characterset. Updated 14 February 1996.

• Linux Firewall HOWTO, by David Rudder <drig@execpc.com>. How to set up a firewall using Linux. Updated 17 July 1995.

• Linux Ftape HOWTO, by Kai Harrekilde-Petersen <khp@dolphinics.no>. Information on ftape drive compatibility with Linux. Updated 22 May 1996 for ftape-2.08.

• Linux GCC HOWTO, by Daniel Barlow <daniel.barlow@linux.org>. How to set up the GNU C compiler and development libraries. Updated 28 February 1996.

• Linux German HOWTO, by Winfried Trümper <truemper@MI.Uni-Koeln.DE>. Information on using Linux with German-specific features. Updated 20 January 1996.

• Linux HAM HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. HOWTO configure amateur radio software for Linux. Updated 30 April 1996.

• Linux HOWTO Index, by Greg Hankins <gregh@sunsite.unc.edu>. Index of HOWTO documents about Linux. Updated 25 August 1996.

• Linux Hardware Compatibility HOWTO, by Tawei Wan <frac@pobox.com>. A list of hardware known to work with Linux. Updated 14 November 1995.

• Linux Hebrew HOWTO, by Yair G. Rajwan <yair@hobbes.jct.ac.il>. How to configure Linux for use with the Hebrew characterset. Updated 12 September 1995.

• Linux INFO-SHEET, by Michael K. Johnson <johnsonm@redhat.com>. Generic introduction to the Linux operating system. Updated 16 July 1996.

• Linux IPX HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to install and configure IPX networking. Updated 1 June 1996.

• Linux ISP Hookup HOWTO, by Egil Kvaleberg <egilk@sn.no>. Basic introduction to hooking up to an ISP. Updated 9 March 1996.

• Linux Installation HOWTO, by Matt Welsh <mdw@sunsite.unc.edu>. How to obtain and install the Linux software. Updated 11 December 1994.

• Linux Italian HOWTO, by Marco “Gaio” Gaiarin <gaio@dei.unipd.it>. How to configure Linux for use with the Italian characterset. Updated 0.6 Beta, 2 Agosto 1996.

• Linux JE HOWTO, by Hiroo Yamagata <hiyori13@interramp.com>. Information on JE, a set of Japanese language extensions for Linux.

• Linux Java HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to get started with Java and HotJava. Updated 29 May 1996.

• Linux Kernel HOWTO, by Brian Ward <ward@blah.tu-graz.ac.at>. Upgrading and compiling the Linux kernel. Updated 1 August 1996.

• Linux Keyboard HOWTO, by Andries Brouwer <aeb@cwi.nl>. Information about the Linux keyboard, console and non-ASCII characters. Updated 8 November 1995.

• Linux META-FAQ, by Michael K. Johnson <johnsonm@redhat.com>. A listing of Linux sources of information. Updated 16 July 1996.

• Linux MGR HOWTO, by Vincent Broman <broman@nosc.mil>. Information on the MGR graphics interface for Linux. Updated 30 May 1996.

• Linux Electronic Mail HOWTO, by Vince Skahan <vince@halcyon.com>. Information on Linux-based mail servers and clients. Updated 29 November 1995.

• Linux Module HOWTO, by Lauri Tischler <ltischler@efore.fi>. How to load modules and a listing of parameters. Updated 20 June 1996.

• Linux NET-2 HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. HOWTO configure TCP/IP networking, SLIP, PLIP, and PPP under Linux. Updated 16 January 1996.

• Linux NIS HOWTO, by Erwin Embsen <erwin@nioz.nl>. Information on using NIS/YP on Linux systems. Updated 24 January 1995.

• Linux News HOWTO, by Vince Skahan <vince@halcyon.com>. Information on USENET news server and client software for Linux. Updated 29 November 1995.

• Linux PCI HOWTO, by Michael Will <Michael.Will@student.uni-tuebingen.de>. Information on PCI-architecture compatibility with Linux. Updated July 1996.

• Linux PCMCIA HOWTO, by Dave Hinds <dhinds@allegro.stanford.edu>. How to install and use PCMCIA Card Services. Updated 28 July 1996.

• Linux PPP HOWTO, by Robert Hart <hartr@hedland.edu.au>. Information on using PPP networking with Linux. Updated 25 August 1996.

• Linux Polish HOWTO, by Sergiusz Pawlowicz <ser@arch.pwr.wroc.pl>. Information on using Linux with Polish-specific features. Updated 24 June 1996.

• Linux Portuguese HOWTO, by Joao Carlos Rodrigues Pereira <jcrp@caravela.di.fc.ul.pt>. How to configure Linux for use with the Portuguese characterset. Updated 14 September 1995.

• Linux Printing HOWTO, by Grant Taylor <gtaylor+pht@picante.com>. HOWTO on printing software for Linux. Updated 25 July 1996.

• Linux Printing Usage HOWTO, by Mark Komarinski <markk@auratek.com>. How to use the printing system for a variety of file types and options. Updated 2 February 1996.

• Linux SCSI HOWTO, by Drew Eckhardt <drew@PoohSticks.ORG>. Information on SCSI driver compatibility with Linux.

• Linux SCSI Programming HOWTO, by Heiko Eissfeldt <heiko@colossus.escape.de>. Information on programming the generic Linux SCSI interface. Updated 7 May 1996.

• Linux SMB HOWTO, by David Wood <dwood@plugged.net.au>. How to use the Session Message Block (SMB) protocol with Linux. Updated 10 August 1996.

• Linux Serial HOWTO, by Greg Hankins <greg.hankins@cc.gatech.edu>. Information on use of serial devices and communications software. Updated 9 October 1995.

• Linux Shadow Password HOWTO, by Michael H. Jackson <mhjack@tscnet.com>. How to obtain, install, and configure shadow passwords. Updated 3 April 1996.

• Linux Sound HOWTO, by Jeff Tranter <Jeff Tranter@Mitel.COM>. Sound hardware and software for the Linux operating system. Updated 10 June 1996.

• Linux Sound Playing HOWTO, by Yoo C. Chung <wacko@power1.snu.ac.kr>. How to play various sound formats under Linux. Updated 18 May 1996.

• Linux Spanish HOWTO, by Gonzalo Garcia Agullo <Gonzalo.Garcia-Agullo@jrc.es>. Information on using Linux with Spanish-specific features. Updated 20 August 1996.

• Linux Term HOWTO, by Patrick Reijnen <patrickr@bart.nl>. HOWTO use the ’term’ communications package on Linux systems. Updated 15 July 1995.

• Linux Tips HOWTO, by Paul Anderson <panderso@ebtech.net>. HOWTO on miscellaneous tips and tricks for Linux. Updated August 1996.

• Linux UMSDOS HOWTO, by Jacques Gelinas <jacques@solucorp.qc.ca>. How to install and use the UMSDOS filesystem. Updated 13 November 1995.

• Linux UPS HOWTO, by Harvey J. Stein <hjstein@math.huji.ac.il>. Information on using a UPS power supply with Linux. Updated 16 April 1995.

• Linux UUCP HOWTO, by Vince Skahan <vince@halcyon.com>. Information on UUCP software for Linux. Updated 29 November 1995.

• Linux XFree86 HOWTO, by Matt Welsh <mdw@sunsite.unc.edu>. How to obtain, install, and configure XFree86 3.1.1 (X11R6). Updated 15 March 1995.

The following mini-HOWTOs are available (plain text only):

• Linux 3 Button Mouse mini-HOWTO, by Geoff Short <grs100@york.ac.uk>. How to configure your mouse to use 3 buttons. Updated 2 July 1996.

• Linux ADSM Backup mini-HOWTO, by Thomas Koenig <Thomas.Koenig@ciw.uni-karlsruhe.de>. How to install and use the ADSM backup program. Updated 3 January 1996.

• Linux Assembly mini-HOWTO, by Francois-Rene Rideau <rideau@ens.fr>. How to program x86 assembly language. Updated 15 June 1996.

• Linux Backup with MSDOS mini-HOWTO, by Christopher Neufeld <neufeld@physics.utoronto.ca>. How to backup Linux machines with MSDOS. Updated 6 February 1996.

• Linux Boca mini-HOWTO, by David H Dennis <david@amazing.cinenet.net>. How to install a Boca 16-port serial card (Boca 2016). Updated 16 February 1995.

• Linux BogoMips mini-HOWTO, by Wim C.A. van Dorst <baron@clifton.hobby.nl>. Information about BogoMips. Updated 16 August 1996.

• Linux Bridge mini-HOWTO, by Chris Cole <chris@polymer.uakron.edu>. How to setup and ethernet bridge. Updated 23 August 1996.

• Linux CD Writer mini-HOWTO, by Matt Cutts <cutts@ms.uky.edu>. How to write CDs. Updated 17 December 1994.

• Linux Colour ls mini-HOWTO, by Thorbjoern Ravn Andersen <ravn@dit.ou.dk>. How to set up the colours with ’ls’. Updated 27 April 1996.

• Linux Consoles mini-HOWTO, by R. Mark Salathiel <RSalathi@nyx.cs.du.edu>. How to set up multiple virtual consoles. Updated 25 December 1994.

• Linux Consoles-Many mini-HOWTO, by Winfried Trümper <truemper@MI.Uni-Koeln.DE>. How to set up many virtual consoles. Updated 17 March 1996.

• Linux DOS2Linux mini-HOWTO, by Guido Gonzato <Guido@ibogfs.cineca.it>. A short guide on getting started with Linux. Updated 26 April 1996.

• Linux Diald mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use ’diald’ to dial an ISP. Updated 3 June 1996.

• Linux Dip+SLiRP+CSLIP mini-HOWTO, by Zenon Fortuna <zenon@netcom.com>. How to use Dip+SLiRP+CSLIP together. Updated 15 October 1995.

• Linux Diskless mini-HOWTO, by Robert Nemkin <buci@math.klte.hu>. How to set up a diskless Linux box. Updated 31 May 1996.

• Linux Dynamic IP Hacks mini-HOWTO, by Michael Driscoll <fenris@lightspeed.net>. Tips and tricks for people with dynamic IP addresses. Updated 22 August 1996.

• Linux GUI Development mini-HOWTO, by Philip Markwalder <pgmarkwa@stud.ee.ethz.ch>. Information about GUI Development Software and Books. Updated 6 June 1995.

• Linux getty-ps mini-HOWTO, by Mark Horton <mah@ka4ybr.com>. How to set up ’getty ps’. Updated 6 September 1994.

• Linux Graphics Tools mini-HOWTO, by Michael J. Hammel <mjhammel@csn.net>. Introduction to graphics tools available for Linux. Updated 15 May 1996.

• Linux Gravis-UltraSound mini-HOWTO, by J-F Mammet <mammet@diva.univ-mlv.fr>. How to install the Gravis UltraSound PnP card. Updated 27 March 1996.

• Linux HTML Validation mini-HOWTO, by Keith M. Corbett <kmc@specialform.com>. How to validate HTML 2.0 documents with nsgmls. Updated 29 October 1995.

• Linux HTTP+Netware mini-HOWTO, by Pramod Karnad <karnadp@mozart.inet.co.th>. How to setup a Linux HTTP server on a Netware LAN. Updated 14 March 1996.

• Linux IO Port Programming mini-HOWTO, by Riku Saikkonen <rjs@spider.compart.fi>. How to use I/O ports in C programs. Updated 26 December 1995.

• Linux IP Alias mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use IP aliasing. Updated 15 August 1996.

• Linux IP Masquerade mini-HOWTO, by Ambrose Au <achau@wwonline.com>. How to use IP masquerading. Updated 17 August 1996.

• Linux Java WorkShop mini-HOWTO, by Mike Gaertner <mg@genyosha.in-chemnitz.de>. How to use Java WorkShop (JWS) for Solaris/intel with Linux. Updated 15 August 1996.

• Linux Jaz Drive mini-HOWTO, by Bob Willmot <bwillmot@cnct.com>. How to use an Iomega Jaz drive with Linux. Updated 5 August 1996.

• Linux Kerneld mini-HOWTO, by Henrik Storner <storner@osiris.ping.dk>. How to use ’kerneld’ (dynamic module loading). Updated 3 June 1996.

• Linux Key Setup mini-HOWTO, by Stephen Lee <sl14@cornell.edu>. How to set up cursor control keys. Updated 16 September 1995.

• Linux Keystroke mini-HOWTO, by Zenon Fortuna <zenon@netcom.com>. How to assign special action to keys. Updated 4 April 1995.

• Linux LBX mini-HOWTO, by Paul D. Smith <psmith@BayNetworks.com>. How to use Low-Bandwidth X (LBX). Updated 23 October 1995.

• Linux LF1000 mini-HOWTO, by Skip Rye <root@brspc 0064.msd.ray.com>. How to use the Panasonic LF1000 optical disk drive. Updated 29 March 1996.

• Linux LILO mini-HOWTO, by Cameron Spitzer <cls@truffula.sj.ca.us>. Examples of typical LILO installations. Updated 4 June 1995.

• Linux Linux+DOS+Win95+OS2 mini-HOWTO, by Mike Harlan <r3mdh@dax.cc.uakron.edu>. How to use Linux and DOS and OS/2 and Win95 together. Updated 6 March 1996.

• Linux Linux+OS2+DOS mini-HOWTO, by Hamish Moffatt <moffatt@yallara.cs.rmit.edu.au>. How to use Linux and OS/2 and DOS together. Updated 20 May 1996.

• Linux Linux+Win95 mini-HOWTO, by Jonathan Katz <jkatz@in.net>. How to use Linux and Windows95 together. Updated 25 June 1996.

• Linux Large Disk mini-HOWTO, Andries Brouwer <aeb@cwi.nl>. How to use disks with > 1024 cylinders. Updated 26 July 1996.

• Linux Locales mini-HOWTO, by Peeter Joot <joot@ecf.toronto.edu>. How to configure Linux to use locales. Updated 6 June 1996.

• Linux Mail Queue mini-HOWTO, by Leif Erlingsson <Leif.Erlingsson@mailbox.swipnet.se>. How to queue remote mail and deliver local mail. Updated 19 September 1995.

• Linux Man Page mini-HOWTO, by Jens Schweikhardt <schweikh@noc.dfn.de>. How to write man pages. Updated April 1996.

• Linux Modeline mini-HOWTO, by Rick Niles <niles@axp745.gsfc.nasa.gov>. How to use ’modeline’. Updated 11 February 1995.

• Linux Multiple Disks Layout mini-HOWTO, by Gjoen Stein <gjoen@nyx.net>. How to layout partitions on multiple disks. Updated 6 August 1996.

• Linux Multiple Ethernet mini-HOWTO, by Don Becker <becker@cesdis.gsfc.nasa.gov>. How to use multiple ethernet cards. Updated 5 August 1995.

• Linux NFS-Root mini-HOWTO, by Andreas Kostyrka <andreas@medman.ag.or.at>. How to set up diskless Linux machines. Updated 14 March 1996.

• Linux Online Support mini-HOWTO, by lilo <TaRDiS@mail.utexas.edu>. Information about the Linux Internet Support Cooperative. Updated 7 May 1996.

• Linux PLIP mini-HOWTO, by Andrea Controzzi <controzz@cli.di.unipi.it>. How to set up PLIP (Parallel Line Interface Protocol). Updated 8 January 1996.

• Linux PPP-over-ISDN mini-HOWTO, by R Marc Phillips <rmarc@netcom.com>. How to set up MLPPP over ISDN. Updated 13 April 1996.

• Linux Process Accounting mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up process accounting. Updated 29 July 1996.

• Linux Proxy ARP mini-HOWTO, by Al Longyear <longyear@netcom.com>. A small treatise on the use of Proxy ARP. Updated 5 December 1994.

• Linux Print2Win mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use SAMBA to print to Windows. Updated 3 June 1994.

• Linux Quota mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up disk quotas. Updated 27 July 1996.

• Linux Reading List, by James H. Haynes <haynes@cats.ucsc.edu>. Interesting books pertaining to Linux subjects. Updated 24 February 1996.

• Linux SLIP+proxyARP mini-HOWTO, by Dave Kennedy <davek@melita.com>. Using SLIP and Proxy ARP to connect LANs to the net. Updated 4 February 1994.

• Linux Stacker mini-HOWTO, by Corey Sweeney <corey@bbs.xnet.com>. How to mount MSDOS Stacked filesystems. Updated 6 September 1994.

• Linux Swap Space mini-HOWTO, by H. Peter Anvin <hpa@yggdrasil.com>. How to share swap space between Linux and Windows. Updated 19 June 1995.

• Linux TIA mini-HOWTO, by Irish <irish@eskimo.com>. How to use ’TIA’ with Linux. Updated 16 December 1995.

• Linux Term Firewall mini-HOWTO, by Barak Pearlmutter <barak.pearlmutter@alumni.cs.cmu.edu>. How to use ’term’ over a firewall. Updated 22 May 1996.

• Linux Tiny News mini-HOWTO, by Kent Lewis <kent@fiona.umsmed.edu>. How to use ’INN’ to make a small news spool. Updated 25 May 1995.

• Linux Token Ring mini-HOWTO, by Mike Eckhoff <meckhoff@zaphod.wayne.esu1.k12.ne.us>. How to use token ring cards. Updated 5 July 1995.

• Linux Upgrade mini-HOWTO, by Greg Louis <glouis@dynamicro.on.ca>. How to upgrade your Linux distribution. Updated 6 June 1996.

• Linux Virtual Web mini-HOWTO, by Dan Pancamo <pancamo@infocom.net>. How to set up a virtual WWW site. Updated 24 November 1995.

• Linux Visual Bell mini-HOWTO, by Alessandro Rubini <rubini@ipvvis.unipv.it>. How to disable audible bells, and enable visual bells. Updated August 1996.

• Linux Win95+Win+Linux mini-HOWTO, by Robert Goodwin <Robert.Goodwin@mcc.ac.uk>. How to get Windows95, Windows 3.x and Linux to work. Updated 4 August 1996.

• Linux WordPerfect mini-HOWTO, by Wade Hampton <tasi029@tmn.com>. How to set up SCO WordPerfect for Linux. Updated 17 May 1995.

• Linux X Notebook mini-HOWTO, by Darin Ernst <darin@castle.net>. How to set up X on notebooks with WD90C24 video chips. Updated 10 August 1995.

• Linux XFree86-XInside mini-HOWTO, by Marco Melgazzi <s64912@athena.polito.it>. How to convert XFree86 to XInside modelines. Updated July 1996.

• Linux Xterminal mini-HOWTO, by Scot W. Stevenson <scot@catzen.gun.de>. How to connect Xterminals to Linux boxen. Updated July 1995.

• Linux ZIP Drive mini-HOWTO, by Grant Guenther <grant@torque.net>. How to use an Iomega ZIP drive with Linux. Updated 15 April 1996.

4 HOWTO Project Developement

4.1 Wanted HOWTOs and mini-HOWTOs

If you are interested in writing about any of these subject, please get in touch with me at

gregh@sunsite.unc.edu.

• Notebook HOWTO or Notebook mini-HOWTO. A guide to notebooks - information about power management, LCD displays, video controlers, pointers to related material (PCMCIA-HOWTO), and most anything pertaining to notebooks.

• If you have an idea for another HOWTO or mini-HOWTO, please get in touch.

4.2 HOWTOs Currently Under Construction

• A Parallel Processing HOWTO (multiprocessor and parallel processing support for Linux) (6 June 1996).

• An Information Services HOWTO is being written (5 March 1996).

• A Microchannel Bus mini-HOWTO is under construction (17 March 1996).

• A Disk HOWTO about disks, partitions, swap space, formatting, filesystems, etc is under construction.

4.3 Writing and Submitting a HOWTO

If you would like to write a Linux HOWTO document, there are a few guidelines that you should follow.

• Get in touch with me first, and let me know you want to work on something. You can reach me

at gregh@sunsite.unc.edu.

• Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• If you are writing a HOWTO, you must use the Linuxdoc-SGML package, available from sunsite.unc.edu:/pub/Linux/utils/text/, to format the HOWTO. This package allows us to produce LaTeX (for DVI and PostScript), plain text, and HTML from a single source document, and was designed by Matt Welsh specifically for the HOWTOs. This also gives all of the HOWTOs a uniform look.

• If you are writing a mini-HOWTO, you can use any form you wish. mini-HOWTOs are meant to be short, free-form HOWTOs on very specific subjects. In general, these should be quick to write and maintain, so it’s not worth bothering with the Linuxdoc-SGML package. This does not mean that the mini-HOWTOs are any less important! Only plain text versions of mini-HOWTOs are distributed.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname), and the current version number and release date of the software.

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an “FAQ” or “Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs as a model. The SGML source to the HOWTOs is available on Linux FTP sites.

• Lastly, be prepared to receive questions and comments about your writing. There are several hundreds of accesses to the HOWTO collection every day from around the world!

After you have written the HOWTO, mail it to me. If you used Linuxdoc-SGML, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTOs on sunsite.unc.edu and posting them to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

Then, all you have to do is send me periodic updates whenever appropriate.

5 Copyright

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at

gregh@sunsite.unc.edu via email.